# Metric Differential Privacy at the User-Level

### v1

@@ 6. Applications of Proposed Mechanisms
6.1. Linear Embedding Queries
Many applications of metric DP assume there is an embedding function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{t}$, which maps an item to its semantic representation in $\mathbb{R}^{t}$ (each of the examples in Sec. 1 have an embedding representation). The metric $d_{\mathcal{X}}$ is then the distance between $\phi(x)$ and $\phi(x^{\prime})$; in this section, we consider the $l_{2}$ distance.
@@ 6.1. Linear Embedding Queries
where $f(y)=Fy$ for a matrix $F\in\mathbb{R}^{d\times t}$ (meaning that $f$ is a linear function). Assume each row $F_{i}$ of $F$ is normalized so that $\|F_{i}\|_{2}\leq 1$. Each coordinate of $f\circ\phi$ is equal to $\mathbb{E}_{x\sim\tilde{K}}[\langle F_{i},\phi(x)\rangle]$. This can be interpreted as the average similarity of each item in $K$ with the vector $F_{i}$. Our analysis will assume that $d<|\mathcal{X}|$ and $d\ll n$, which is usually the case in practice. Note that we may write $q_{f\circ\phi}$ as $F\Phi\tilde{K}$, where $\Phi\in\mathbb{R}^{t\times\mathcal{X}}$ is the collection of embedding vectors in $\mathcal{X}$.
6.1.1. Local Model
Existing user-level DP solutions ask each user $U_{i}$ to privately release the query $\hat{q}_{i}=q_{f\circ\phi}(\tilde{K}_{i})$. The aggregator computes the average $\hat{q}=\frac{1}{n}\sum_{i=1}^{n}\hat{q}_{i}$. The current best solutions have the following error guarantee (Duchi et al., 2013; Bassily, 2019):
@@ 6.1.1. Local Model
(From Proposition 3 in (Duchi et al., 2013)) There exists an $(\varepsilon,0)$-bounded user-level DP in the local model algorithm which produces an estimate $\hat{q}$ such that, for all $\tilde{K}$,
@@
On the other hand, for $d_{\textsf{EM}}$-DP, by Thm. 4.1, we know that $\Delta_{\mathsf{EM}}(q_{f\circ\phi})$ is at most the Lipschitz constant of $f\circ\phi$ given by:
@@
Remarks. We use the Gaussian mechanism because it performs better under the $\ell_{2}$ error than the pure $(\alpha,0)$-bounded local $d_{\textsf{EM}}$ DP illustrated in Alg. 1. However, this forces us to use $\delta>0$. We leave it as an interesting open question whether similar error can be achieved with $\delta=0$. Compared to Lemma 6.1, the above bound differs by a factor of $\frac{\varepsilon}{\alpha}$ (and small $\ln\frac{1}{\delta}$ terms)—when $\alpha=\varepsilon$, we know that $d_{\textsf{EM}}$-DP provides better privacy. When $\varepsilon\ll\alpha$, that PrivEMDItemWise offers lower error than Lemma 6.1.
6.1.2. Central Model
In the central model, linear query release has been extensively studied, and optimal algorithms under item-level DP are known (Hardt and Talwar, 2010; Bhaskara et al., 2012; Nikolov et al., 2013). These algorithms can be easily adapted to user-level DP, which will provide the following guarantee:
@@ 6.1.2. Central Model
(From Thm. 1.3 in (Hardt and Talwar, 2010)) There exists an $(\varepsilon,0)$-bounded user-level DP algorithm in the central model which produces an estimate $\hat{q}$ such that, for all $\tilde{K}$,
@@
6.2. Frequency Estimation
Here, we evaluate the error of PrivEMDItemWise for private frequency estimation, where the goal is to obtain a private estimate $\tilde{F}$ of the (normalized) histogram $\tilde{K}$. This problem has been extensively studied in privacy (Hay et al., 2009; Xu et al., 2013; Suresh, 2019; Kairouz et al., 2016; Acharya et al., 2018; Chen et al., 2020; Acharya et al., 2023); the high-level goal is to minimize the $\ell_{p}$ distance between $\tilde{F}$ and $\tilde{K}$. However when the data domain is a general metric space $\mathcal{X}$, not all $\ell_{p}$ perturbations to $\tilde{K}$ are the same. Thus, we will measure similarity between $\tilde{K},\tilde{F}$ with $d_{\textsf{EM}}(\tilde{K},\tilde{F})$, as we do in our privacy definition.
To ease the analysis while still demonstrating the effectiveness of our mechanisms, we fix $\mathcal{X}$ to be the following “clustered” metric space. Let $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, where =$\mathcal{B}=\{b_{1},\ldots,b_{s}\}$, $\mathcal{C}=\{c_{1},\ldots,c_{t}\}$ and $s\cdot t=k$. For some $r<\frac{1}{2}$, the distance is given by the following:
@@ 6.2. Frequency Estimation
6.2.1. Algorithms in the Local Model
At a high level, in the local model each user applies a private mechanism $\mathcal{A}:\mathcal{X}\rightarrow\mathcal{Y}$ (with $\mathcal{Y}$ discrete and $|\mathcal{Y}|\geq|\mathcal{X}|$) to each sample and releases it. The central server forms an aggregate vector $v\in\mathbb{R}^{\mathcal{Y}}$. Let $A\in\mathbb{R}^{\mathcal{X}\times\mathcal{Y}}$ denote the transition probability matrix of $\mathcal{A}$; we have by linearity of expectation that $\mathbb{E}[v]=\tilde{K}A$. Assuming that $A$ has a right inverse $B$, the central server returns the estimate $\tilde{L}=vB$, which is unbiased. All previous work in distribution estimation under local DP can be expressed in this way (Kairouz et al., 2016; Acharya et al., 2018; Chen et al., 2020; Acharya et al., 2023). We summarize this in Alg. 4.
@@ 6.2.1. Algorithms in the Local Model
$\tilde{F}=vB$;
return $\tilde{F}$
@@
The state-of-the-art approach for frequency estimation is the Hadamard response (Acharya et al., 2018; Chen et al., 2020), which is based off of the Hadamard matrices (which form a robust encoding of $\mathcal{X}$). Specifically, the matrix $A$ is given by $q_{1}\textbf{1}+q_{2}H$, where $H$ is a Hadamard matrix and $q_{1},q_{2}$ are constants chosen so that $A$ is normalized and that each element is proportional to either $e^{\varepsilon}$ or $1$. This mechanism has the following utility:
@@
(From Thm. 3.1 in (Chen et al., 2020)) There exists a mechanism $\mathcal{A}$ such that FreqEstLocal satisfies $(\varepsilon,\delta)$-bounded user-level DP and returns an estimator $\tilde{F}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{K},\tilde{F})]\leq O\left(\sqrt{\tfrac{k}{mn}}+\sqrt{\tfrac{k^{2}\ln\frac{m}{\delta}}{n\varepsilon^{2}}}\right).$
Remarks. In order to adapt the Hadamard response to the user-level setting, we suppose each user applies $\mathcal{A}$ to each sample with privacy budget $\frac{\varepsilon}{\sqrt{m\ln(m/\delta)}}$, and $(\varepsilon,\delta)$-user level DP follows from composition (Kairouz et al., 2015). The term $\sqrt{\frac{k}{mn}}$ is a sampling error which does not depend on $\varepsilon$, and the second $\sqrt{\frac{k^{2}\ln(m/\delta)}{n\varepsilon^{2}}}$ term is the cost of privacy. The cost of privacy usually dominates, and furthermore its dependence on $m$ is not significant. This is because $m$ reduces both the effect of each sample on the final estimate, and the privacy budget per sample, countervailing itself.
Under $d_{\textsf{EM}}$-DP, we can use a transition probability matrix $A$ that is less noisy. Specifically, each user may apply $\mathcal{A}$ to their dataset using PrivEMDItemWise, and by Thm. 4.3, $\mathcal{A}$ needs to satisfy $(O(\frac{\alpha}{\sqrt{m\ln(me^{\alpha}/\delta)}}),0)$ $d_{\mathcal{X}}$-DP. Note that for our choice of $\alpha$ and $\mathcal{X}$, this is often a less restrictive requirement than ($\frac{\varepsilon}{\sqrt{m\ln(m/\delta)}},0)$-DP) since $\alpha<\varepsilon$. We first derive an error bound on FreqEstLocal in terms of $A$ (specifically its right inverse), which we will then optimize later.
@@
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$ and any mechanism $\mathcal{A}$ satisfying ($\alpha_{0},0)$ $d_{\mathcal{X}}$-DP where $\alpha_{0}=O(\frac{\alpha}{\sqrt{m\ln(me^{\alpha}/\delta)}})$ ($\alpha_{0}$ is specifically defined in Thm. 4.4), FreqEstLocal satisfies $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP in the local model and returns an estimator $\tilde{F}$ such that
(8) $\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{F},\tilde{K})]\leq r\sqrt{\frac{st(\|B^{T}\|_{1\rightarrow 2}^{2}-1)}{mn}}+\sqrt{\frac{s(\|P^{T}B^{T}\|_{1\rightarrow 2}^{2}-1)}{mn}},$
@@
Remarks. The first term in the RHS of (8) is the cost of equalizing mass between clusters, and the second term is the cost of equalizing the mass across clusters (since the matrix $P$ essentially projects $\mathcal{A}$ to act between clusters). For small $r$, the first term approaches 00, and the latter term may also approach 00 because $\mathcal{A}$ will not often map a point outside its cluster under $d_{\mathcal{X}}$-DP (and thus, $\|P^{T}B^{T}\|^{2}_{1\rightarrow 2}-1\rightarrow 0$).
Proof Sketch. Our bound forms a transportation plan between $\tilde{F}$ and $\tilde{K}$ by first mapping the mass within each cluster arbitrarily, which incurs at most $r\|\tilde{F}-\tilde{K}\|_{1}$ cost, and then equalizing the mass between clusters, which incurs at most $\|\tilde{F}-\tilde{K}\|_{1}$ cost. Both of the error terms can then be bounded by viewing $\tilde{F}-\tilde{K}$ as the sum of $mn$ independent variables drawn from a Dirichlet distribution with mean 00, and applying a standard variance analysis.∎
We apply Thm. 6.5 with $\mathcal{A}$ being a generalization of $k$-randomized response (Kairouz et al., 2016) which is adapted to $d_{\mathcal{X}}$-DP. Specifically, $\textsf{GKRR}_{\alpha_{0}}$ has probabilities given by, for each $(b,c)\in\mathcal{X}$,
@@
$\Pr[\textsf{GKRR}(b,c))=(b^{\prime},c^{\prime})]\propto 1~{}~{}~{}~{}~{}~{}\forall b^{\prime}\neq b,c.$
@@
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, FreqEstLocal with the mechanism $\mathcal{A}=\textsf{GKRR}_{\alpha_{0}}$ satisfies $(\alpha,\delta)$-$d_{\textsf{EM}}$ DP in the local model and returns an estimator $\tilde{F}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{F},\tilde{K})]\leq r\sqrt{\frac{st^{3}}{mn}}\left(\frac{e^{\alpha_{0}}+s}{e^{\alpha_{0}}-e^{(1-r)\alpha_{0}}}\right)\\ +\sqrt{\frac{s^{2}t^{2}}{mn}}\left(\frac{\sqrt{s+2(e^{\alpha_{0}}-1)}}{e^{\alpha_{0}}+(t-1)e^{(1-r)\alpha_{0}}-t}\right),$
@@
Remarks: Specifically, for our choice of $\alpha=\frac{\varepsilon}{r}$, we have
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{K},\tilde{F})]\leq 4\sqrt{\tfrac{k^{3}}{mn}}+64\tfrac{\sqrt{k^{3}}}{\alpha\sqrt{n}}\sqrt{\ln(4m\exp(\alpha)/\delta)}.$
Similar to Lemma 6.4, the $\sqrt{\frac{k^{3}}{mn}}$ term is the cost of sampling. The $r\frac{\sqrt{k^{3}}}{\varepsilon\sqrt{n}}$ term is the cost of privacy, and it dominates when $\alpha\leq\sqrt{m}$. We will compare Thm. 6.6 with Lemma 6.4 when $k,\varepsilon,\alpha<\sqrt{m}$—then the cost of privacy dominates. Specifically, the cost of Lemma 6.4 is $O(\sqrt{\frac{k^{2}\ln(m/\delta)}{n\varepsilon^{2}}})$, and the cost of Thm. 6.6 is $O(\sqrt{\frac{k^{3}}{\alpha^{2}n}\max\{\ln(\frac{m}{\delta}),\alpha\}})$. Given $\varepsilon$, the error will be smaller if
@@
i.e. if there is a gap between $\alpha,\varepsilon$ of size at least $\sqrt{k}$. This is possible if $k\ll\frac{1}{r}$, and for these instances $d_{\textsf{EM}}$ DP offers better utility than user-level DP. In Thm. 6.6, the super-linear factor of $k^{3/2}$ comes from the fact that the $k$-RR is suboptimal in terms of $k$ (Acharya et al., 2018).
6.2.2. Algorithms in the Central Model
The Laplace mechanism has been shown to be optimal for many instances of frequency estimation (Dwork et al., 2014). To attain user-level privacy, the baseline Laplace mechanism releases, for each $x\in\mathcal{X}$, the values $F_{x}=\tilde{K}_{G}(x)+Y$, where $Y\sim Lap(\frac{1}{n\varepsilon})$. The distribution function $\tilde{F}$ is then the normalization of $\langle F_{x}:x\in\mathcal{X}\rangle$. This gives us the following guarantees.

### v2

@@ 6 Applications of Proposed Mechanisms
6.1 Linear Embedding Queries
Many applications of metric DP assume there is an embedding function $\phi:\mathcal{X}\rightarrow\mathbb{R}^{t}$, which maps an item to its semantic representation in $\mathbb{R}^{t}$ (each of the examples in Section 1 have an embedding representation). The metric $d_{\mathcal{X}}$ is then the distance between $\phi(x)$ and $\phi(x^{\prime})$; in this section, we consider the $l_{2}$ distance.
@@ 6.1 Linear Embedding Queries
where $f(y)=Fy$ for a matrix $F\in\mathbb{R}^{d\times t}$ (meaning that $f$ is a linear function). Assume each row $F_{i}$ of $F$ is normalized so that $\|F_{i}\|_{2}\leq 1$. Each coordinate of $f\circ\phi$ is equal to $\mathbb{E}_{x\sim\tilde{K}}[\langle F_{i},\phi(x)\rangle]$. Thus, we may view each coordinate of a linear embedding query as a similarity query in the embedding space with query point $F_{i}$. Our analysis will assume that $d<|\mathcal{X}|$ and $d\ll n$, which is usually the case in practice. Note that we may write $q_{f\circ\phi}$ as $F\Phi\tilde{K}$, where $\Phi\in\mathbb{R}^{t\times\mathcal{X}}$ is the collection of embedding vectors in $\mathcal{X}$.
6.1.1 Local Model
Existing user-level DP mechanisms ask user at index $i$ to privately release the query $\hat{q}_{i}=q_{f\circ\phi}(\tilde{K}_{i})$. The aggregator computes the average $\hat{q}=\frac{1}{n}\sum_{i=1}^{n}\hat{q}_{i}$. The current best solutions have the following error guarantee (Duchi et al., 2013; Bassily, 2019) [^4] No algorithms are known which satisfy $(\varepsilon,\delta)$-DP and have better error than the $(\varepsilon,0)$-DP algorithm shown.:
@@ 6.1.1 Local Model
(From Proposition 3 in Duchi et al. (2013)) There exists an $(\varepsilon,0)$-bounded user-level DP in the local model algorithm which produces an estimate $\hat{q}$ such that, for all $\tilde{K}$,
@@
On the other hand, for $d_{\textsf{EM}}$-DP, by Theorem 4.1, we know that $\Delta_{\mathsf{EM}}(q_{f\circ\phi})$ is at most the Lipschitz constant of $f\circ\phi$ given by:
@@
Remarks. We use the Gaussian mechanism because it performs better under the $\ell_{2}$ error than the pure $(\alpha,0)$-bounded local $d_{\textsf{EM}}$ DP illustrated in Algorithm 1. However, this forces us to use $\delta>0$. Compared to Lemma 6.1, the above bound differs by a factor of $\frac{\varepsilon}{\alpha}$ (and small $\ln\frac{1}{\delta}$ terms)—when $\alpha=\varepsilon$, we know that $d_{\textsf{EM}}$-DP provides better privacy. When $\varepsilon\ll\alpha$, that PrivEMDItemWise offers lower error than Lemma 6.1.
6.1.2 Central Model
In the central model, linear query release has been extensively studied, and optimal algorithms under item-level DP are known (Hardt and Talwar, 2010; Bhaskara et al., 2012; Nikolov et al., 2013). These algorithms can be easily adapted to user-level DP, which will provide the following guarantee [^5] For simplicity, we do not state or compare to the exact instance-optimal upper bounds known for linear queries. Instead, the upper bound in Lemma 6.3 is the optimal one for random linear queries Hardt and Talwar (2010). Like in the local case, there is no separation between $(\varepsilon,\delta)$-DP and $(\varepsilon,0)$-DP for this problem.:
@@ 6.1.2 Central Model
(From Theorem 1.3 in Hardt and Talwar (2010)) There exists an $(\varepsilon,0)$-bounded user-level DP algorithm in the central model which produces an estimate $\hat{q}$ such that, for all $\tilde{K}$,
@@
6.2 Frequency Estimation
Here, we evaluate the error of PrivEMDItemWise for private frequency estimation, where the goal is to obtain a private estimate $\tilde{H}$ of the (normalized) histogram $\tilde{K}$. This problem has been extensively studied in privacy Hay et al. (2009); Xu et al. (2013); Suresh (2019); Kairouz et al. (2016); Acharya et al. (2018); Chen et al. (2020); Acharya et al. (2023); the high-level goal is to minimize the $\ell_{p}$ distance between $\tilde{H}$ and $\tilde{K}$. However when the data domain is a general metric space $\mathcal{X}$, not all $\ell_{p}$ perturbations to $\tilde{K}$ are the same. Therefore, we measure the similarity between $\tilde{K},\tilde{H}$ via $d_{\textsf{EM}}(\tilde{K},\tilde{H})$, as we do in our privacy definition.
To simplify the analysis while still demonstrating the effectiveness of our mechanisms, we fix $\mathcal{X}$ to be the following “clustered” metric space. Let $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, where =$\mathcal{B}=\{b_{1},\ldots,b_{s}\}$, $\mathcal{C}=\{c_{1},\ldots,c_{t}\}$ and $s\cdot t=k$. For some $r<\frac{1}{2}$, the distance is given by the following:
@@ 6.2 Frequency Estimation
6.2.1 Algorithms in the Local Model
At a high level, in the local model each user applies a private mechanism $\mathcal{A}:\mathcal{X}\rightarrow\mathcal{Y}$ (with $\mathcal{Y}$ discrete and $|\mathcal{Y}|\geq|\mathcal{X}|$) to each sample and releases it. The central server forms an aggregate vector $v\in\mathbb{R}^{\mathcal{Y}}$. Let $A\in\mathbb{R}^{\mathcal{X}\times\mathcal{Y}}$ denote the transition probability matrix of $\mathcal{A}$; we have by linearity of expectation that $\mathbb{E}[v]=\tilde{K}A$. Assuming that $A$ has a right inverse $B$, the central server returns the estimate $\tilde{H}=vB$, which is unbiased. All previous work in distribution estimation under local DP can be expressed in this way (Kairouz et al., 2016; Acharya et al., 2018; Chen et al., 2020; Acharya et al., 2023). We summarize this in Algorithm 4.
@@ 6.2.1 Algorithms in the Local Model
$\tilde{H}=vB$;
return $\tilde{H}$
@@
The state-of-the-art approach for frequency estimation is the Hadamard response Acharya et al. (2018); Chen et al. (2020), which is based off of the Hadamard matrices (which form a robust encoding of $\mathcal{X}$). Specifically, the matrix $A$ is given by $q_{1}\textbf{1}+q_{2}H$, where $H$ is a Hadamard matrix and $q_{1},q_{2}$ are constants chosen so that $A$ is normalized and that each element is proportional to either $e^{\varepsilon}$ or $1$. This mechanism has the following utility:
@@
(From Theorem 3.1 in Chen et al. (2020)) There exists a mechanism $\mathcal{A}$ such that FreqEstLocal satisfies $(\varepsilon,\delta)$-bounded user-level DP and returns an estimator $\tilde{H}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{K},\tilde{H})]\leq O\left(\sqrt{\tfrac{k}{mn}}+\sqrt{\tfrac{k^{2}\ln(m/\delta))}{n\varepsilon^{2}}}\right).$
Remarks. In order to adapt the Hadamard response to the user-level setting, we suppose each user applies $\mathcal{A}$ to each sample with privacy budget $\frac{\varepsilon}{\sqrt{m\ln(m/\delta)}}$, and $(\varepsilon,\delta)$-user level DP follows from composition Kairouz et al. (2015). The term $\sqrt{\frac{k}{mn}}$ is the sampling error which does not depend on $\varepsilon$, and the second $\sqrt{\frac{k^{2}\ln(m/\delta)}{n\varepsilon^{2}}}$ term is the cost of privacy. The cost of privacy usually dominates, and furthermore its dependence on $m$ is not significant. This is because $m$ reduces both the effect of each sample on the final estimate, and the privacy budget per sample, countervailing itself.
With $d_{\textsf{EM}}$-DP under our chosen metric, we can use a transition probability matrix $A$ that is less noisy. This comes from the fact that only items in the same cluster need a strong privacy guarantee, whereas traditional user-level DP would require all points to have this level of privacy guarantee. To provide formal guarantees, we first derive an error bound on FreqEstLocal in terms of $A$ (specifically its right inverse), which we will then optimize later.
@@
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$ and any mechanism $\mathcal{A}$ satisfying ($\alpha_{0},0)$ $d_{\mathcal{X}}$-DP where $\alpha_{0}=O(\frac{\alpha}{\sqrt{m\ln(me^{\alpha}/\delta)}})$ ($\alpha_{0}$ is specifically defined in Theorem 4.4), FreqEstLocal is $(\alpha,\delta)$-bounded $d_{\textsf{EM}}$-DP in the local model and returns an estimator $\tilde{H}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{H},\tilde{K})]\leq r\sqrt{\frac{st(\|B^{T}\|_{1\rightarrow 2}^{2}-1)}{mn}}+\sqrt{\frac{s(\|P^{T}B^{T}\|_{1\rightarrow 2}^{2}-1)}{mn}},$ (8)
@@
Remarks. The first term in the RHS of Eq. (8) is the cost of equalizing the mass between clusters, and the second term is the cost of equalizing the mass across clusters (since the matrix $P$ essentially projects $\mathcal{A}$ to act between clusters). For small $r$, the first term approaches 00, and the latter term may also approach 00 because $\mathcal{A}$ will not often map a point outside its cluster under $d_{\mathcal{X}}$-DP (and thus, $\|P^{T}B^{T}\|^{2}_{1\rightarrow 2}-1\rightarrow 0$).
Proof Sketch. Observe that we may upper bound $d_{\textsf{EM}}(\tilde{H},\tilde{K})$ with any transportation plan between $\tilde{H},\tilde{K}$. We will use the following one: first map the probability masses in each cluster so that they match, putting extra mass in an arbitrary point. This incurs at most $r\|\tilde{H}-\tilde{K}\|_{1}$ cost, since the intra-cluster distance is at most $r$. Next, equalizing the mass between clusters, which incurs at most $\|P(\tilde{H}-\tilde{K})\|_{1}$ cost, where $P$ is the given matrix which can be viewed as the linear operator that adds the mass within each cluster together. Both of the error terms can then be bounded by viewing $\tilde{H}-\tilde{K}$ as the sum of $mn$ independent variables drawn from a Dirichlet distribution with mean 00, and applying a variance analysis.∎
Now, the task is to pick a mechanism $\mathcal{A}$ satisfying $d_{\textsf{EM}}$-DP, which minimizes the error in (8). The constraint of $d_{\textsf{EM}}$-DP is quite different from standard DP, and permits novel mechanism design. We use a natural generalization of $k$-randomized response Kairouz et al. (2016), adapted to $d_{\mathcal{X}}$-DP. Specifically, $\textsf{GKRR}_{\alpha_{0}}$ has probabilities given by, for each $(b,c)\in\mathcal{X}$,
@@
$\Pr[\textsf{GKRR}(b,c))=(b^{\prime},c^{\prime})]\propto 1~{}~{}~{}~{}~{}~{}\forall b^{\prime}\neq b,c^{\prime}.$
@@
For the metric space $\mathcal{X}=\mathcal{B}\times\mathcal{C}$, FreqEstLocal with the mechanism $\mathcal{A}=\textsf{GKRR}_{\alpha_{0}}$ satisfies $(\alpha,\delta)$-$d_{\textsf{EM}}$ DP in the local model and returns an estimator $\tilde{H}$ such that
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{H},\tilde{K})]\leq r\sqrt{\frac{st^{3}}{mn}}\left(\frac{e^{\alpha_{0}}+s}{e^{\alpha_{0}}-e^{(1-r)\alpha_{0}}}\right)+\sqrt{\frac{s^{2}t^{2}}{mn}}\left(\frac{\sqrt{s+2(e^{\alpha_{0}}-1)}}{e^{\alpha_{0}}+(t-1)e^{(1-r)\alpha_{0}}-t}\right),$ (9)
@@
Remarks. Specifically, for our choice of $\alpha=\frac{\varepsilon}{r}$, we have
$\max_{K}\mathbb{E}[d_{\textsf{EM}}(\tilde{K},\tilde{H})]\leq 4\sqrt{\tfrac{k^{3}}{mn}}+64\tfrac{\sqrt{k^{3}}}{\alpha\sqrt{n}}\sqrt{\ln(4m\exp(\alpha)/\delta)}.$
Similar to Lemma 6.4, the $\sqrt{\frac{k^{3}}{mn}}$ term is the cost of sampling. The $r\frac{\sqrt{k^{3}}}{\varepsilon\sqrt{n}}$ term is the cost of privacy, and it dominates when $\alpha\leq\sqrt{m}$. We will compare Theorem 6.6 with Lemma 6.4 when $k,\varepsilon,\alpha<\sqrt{m}$—then the cost of privacy dominates. Specifically, the cost of Lemma 6.4 is $O(\sqrt{\frac{k^{2}\ln(m/\delta)}{n\varepsilon^{2}}})$, and the cost of Theorem 6.6 is $O(\sqrt{\frac{k^{3}}{\alpha^{2}n}\max\{\ln(\frac{m}{\delta}),\alpha\}})$. Given $\varepsilon$, the error will be smaller if
@@
i.e. if there is a gap between $\alpha,\varepsilon$ of size at least $\sqrt{k}$. This is possible if $k\ll\frac{1}{r}$, and for these instances $d_{\textsf{EM}}$ DP offers better utility than user-level DP. In Theorem 6.6, the super-linear factor of $k^{3/2}$ comes from the fact that the $k$-RR is suboptimal in terms of $k$ (Acharya et al., 2018).
6.2.2 Algorithms in the Central Model
The Laplace mechanism has been shown to be optimal for many instances of frequency estimation (Dwork et al., 2014). To attain user-level privacy, the baseline Laplace mechanism releases, for each $x\in\mathcal{X}$, the values $F_{x}=\tilde{K}_{G}(x)+Y$, where $Y\sim Lap(\frac{1}{n\varepsilon})$. The distribution function $\tilde{H}$ is then the normalization of $\langle F_{x}:x\in\mathcal{X}\rangle$. This gives us the following guarantee.